The present sum of father & son is 50.When son reached at father present age,the sum of father & son age will be 102.Now calculate father & son age?
Answer by jim_thompson5910(35256)

f = current age of father
s = current age of son

Their current ages add to 50
f+s = 50
Call this equation 1.

t years later,
(f+t) + (s+t) = 102

t = difference in father and son's age
t = f - s

Plug this in
(f+t) + (s+t) = 102
(f+f - s) + (s+f - s) = 102
3f - s = 102
Call this equation 2.


We have the system
f + s = 50
3f - s = 102
which are equation 1 and equation 2 respectively

Add the equations straight down to get
f + s = 50
3f - s = 50
-----------
4f + 0s = 152

Since 4f = 152, we know f = 38 after we divide both sides by 4.

Therefore, s = 50-f = 50-38 = 12

Summary:

The father's current age is 38
The son's current age is 12


Note how the difference in their age is 38-12 = 26
If 26 years pass by, then
father's future age = 38+26 = 64
son's future age = 12+26 = 38
those future ages add up to 64+38 = 102